OpenStudy (anonymous):

Will medal!! What is the dimensions of an octago nwith the area of 154cm?

OpenStudy (stamp):

Referencing your textbook, lecture, online, or notes, what is the formula for the area of an octagon?

OpenStudy (anonymous):

A=2(1+2)a2

OpenStudy (anonymous):

@stamp

OpenStudy (anonymous):

are there any options?

OpenStudy (anonymous):

Sadly no

OpenStudy (stamp):

\[A=2(1+\sqrt{2})a^2\] A is area, and a is the side length. We don't need options! this is math

OpenStudy (stamp):

They told us already A = 195. Substituting 195 for A into our formula\[195=2(1+\sqrt{2})a^2\]

OpenStudy (stamp):

Are you able to get the a-squared by itself?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks i think i figured it out

OpenStudy (stamp):

\[\frac{195}{2(1+\sqrt{2})}=a^2\]

OpenStudy (stamp):

\[\sqrt{\frac{195}{2(1+\sqrt{2})}}=a\]